Some people recommend that the United States impose restrictions on imports in order to reduce its current account deficit.   When is this approach likely to be effective?   Explain your answer in terms of the national income identity CA = (Sp – I) + (T – G).
Answer #1

What is a Current Account Deficit

The current account deficit is a measurement of a country’s trade where the value of the goods and services it imports exceeds the value of the goods and services it exports. The current account includes net income, such as interest and dividends, and transfers, such as foreign aid, although these components make up only a small percentage of the total current account. The current account represents a country’s foreign transactions and, like the capital account is a component of a country’s balance of payments.

The components of current account are

  1. Trade balance The exports and imports of goods and services. Exports lead to payments from the rest of the world, imports to payments to the rest of the world;

    2. Investment income Income on the residents’ holdings of foreign assets, and payments to foreign residents’ holdings of the country’s assets;

    3. Net transfer received The net value of a country’s payments on giving and receiving foreign aid.

The components of the balance of payments are

1. The current account;

2. The capital account, or the net capital flows In the balance of payments, a summary of a country’s asset transactions with the rest of the world;

3. Statistical discrepancy In principle, the net capital flows should be equal to the current account deficit. However, in practice, the difference between these two is not zero, which mostly comes from the imperfect economic data. It is called Statistical discrepancy

                   

                    Higher barriers to imports may have little or no impact upon private savings, investment, and the budget deficit. If there were no effect on these variables, then the current account would not improve with the imposition of tariffs or quotas. It is possible to tell stories in which the effect on the current account goes either way. For example, investment could rise in industries protected by the tariff, worsening the current account. (Indeed, tariffs are sometimes justified by the alleged need to give ailing industries a chance to modernize their plant and equipment.) On the other hand, investment might fall in industries that face a higher cost of imported intermediate goods as a result of the tariff. In general, permanent and temporary tariffs have different effects. The point of the question is that a prediction of the manner in which policies affect the current account requires a general-equilibrium, macroeconomic analysis.

                 

               One country’s GDP Y = C + I + G + X − IM and CA = X − IM + S P, we get Y = C +I +G +CA−S P. The identity of the residents’ income Y +S P = C +T +S in which S is the saving. Combine these two to get I +G +CA = T + S , i.e. CA = T −G + S − I in which G − T is called the government budget deficit. Therefore to reduce a current account deficit, one country can increase its private saving, reduce domestic investment, or cut its government budget deficit.